Calculate the concentrations in g dm-3 of the ions in the following solutions:

a.chlorine ion in 0.3M sodium chloride solution

b.NO3 ions in 0.75M trioxonitrate (v) acid

c.SO3 ions in 3M potassium tetraoxosulphate (iv) solution

1
Expert's answer
2021-01-25T04:33:27-0500

a. T(Cl-)=m(Cl-)/V(solution)

m(Cl-)=n*M=0.3*35.5=10.65 g

So

T=10.65/V; g/dm3

b. T(NO3-)=m(NO3-)/V(solution)

m(Cl-)=n*M=0.75*62=46.5 g

So

T=46.5/V; g/dm3

c. T(SO32-)=m(SO32-)/V(solution)

m(SO3-)=n*M=3*80=240 g

So

T=240/V; g/dm3


It is not enough data to solve this exercise. I mean the volume of solutions.